# MERA Code: A Unified Framework for Evaluating Code Generation Across Tasks

### v2

@@ 4.2 Results
GPT-4.1 0.377 0.382 0.144 0.418 0.416 0.443 0.45 0.344 0.555 0.096 0.162 0.297 0.66
GPT-4o 0.377 0.381 0.149 0.324 0.399 0.529 0.537 0.37 0.479 0.078 0.153 0.275 0.666
Gemini 2.5 flash 0.356 0.427 0.12 0.388 0.386 0.61 0.604 0.211 0.496 0.064 0.174 0.217 0.404
DeepSeek-Coder-V2-Inst 0.347 0.36 0.149 0.363 0.383 0.433 0.392 0.269 0.494 0.06 0.153 0.2 0.714
GigaChat 2 Max 0.346 0.365 0.106 0.332 0.342 0.537 0.53 0.137 0.425 0.062 0.175 0.294 0.68
Qwen2.5-Coder-32B-Inst 0.296 0.306 0.111 0.323 0.383 0.311 0.289 0.22 0.466 0.065 0.168 0.213 0.519
GigaCode 1.4 0.289 0.166 0.135 0.322 0.352 0.357 0.305 0.313 0.027 0.064 0.189 0.276 0.676
Qwen2.5-72B-Inst 0.285 0.254 0.144 0.362 0.349 0.174 0.157 0.189 0.481 0.048 0.128 0.252 0.702
Seed-Coder-8B-Inst 0.268 0.345 0.106 0.106 0.305 0.317 0.21 0.264 0.655 0.035 0.128 0.237 0.403
Yi-Coder-9B-Chat 0.203 0.181 0.135 0.067 0.252 0.35 0.173 0.229 0.145 0.016 0.138 0.192 0.364
Mixtral-8x22B-Inst 0.179 0.028 0.106 0.18 0.366 0.0 0.0 0.229 0.027 0.016 0.159 0.152 0.597
Vikhr-YandexGPT-5-Lite-8B 0.168 0.187 0.091 0.032 0.201 0.035 0.024 0.123 0.407 0.022 0.106 0.138 0.464
@@
Table 2: MERA Code benchmark results. The private tasks are CodeLinterEval, ruCodeEval and ruCodeReviewer.
Table 2 shows the results of the baselines evaluation. Gemini 2.5 Flash (0.356), GPT-4o (0.377) and GPT-4.1 (0.377) demonstrate superior overall performance across eleven coding tasks, particularly excelling in multi-language documentation (StRuCom), Python/Java completions (RealCode/RealCodeJava/ruHumanEval/ruCodeEval), and unit-test generation. They are closely trailed by DeepSeek-Coder-V2 (0.347) and GigaChat-2-Max (0.346), with the former showing exceptional capability in predicting code compilation success (CodeCorrectness: 0.714). Vikhr-YandexGPT-5-Lite lags significantly (0.168), showing near-zero performance in algorithmic and code completion tasks. All models exhibit pronounced weaknesses in generating unit-tests on Python (UnitTests) and automated comments generation (ruCodeReviewer), where maximum scores remain below 0.1 despite task apparent simplicity.
Specialized models show targeted strengths: DeepSeek-Coder-V2 excels in compilation prediction (0.714 vs GPT-4o’s 0.666), while Seed-Coder-8B dominates error correction (0.65 for CodeLinterEval - about 35% better than the average result). Mixtral-8x22B outperforms GPT-4o on multilingual unit-tests generation (+4%). Qwen2.5-72B demonstrates worse results compared to Qwen2.5-Coder-32B (-4-20% on half of the tasks) differing the most on comments generation (-25%), predicting code executability (+28%), code generation on C/C++ (+23%) and algorithmic tasks (-45%). GigaChat-2-Max shows strength in complex Python completions (RealCode: 0.332, ruCodeEval: 0.537), and GPT-4.1 maintains consistent documentation quality (StRuCom: 0.297). GigaCode 1.4 achieves the best score in the generation of multilingual unit-tests (0.188). However, inconsistent JUnit5 test generation highlights critical gaps. These results affirm that while versatile architectures like GPT-4 excel broadly, targeted fine-tuning yields niche advantages—though reasoning and debugging capabilities remain key challenges across all models. Extended results could be found in Appendix C.
5 Conclusion
@@
In this work, we present the MERA Code, a comprehensive evaluation tool designed for coding tasks and programming languages using Russian prompts across eight different languages. MERA provides an open-source evaluation toolkit, a structured scoring system, and a public leaderboard, enabling standardized assessments of coding skills while addressing limitations in non-English contexts. This framework enhances our understanding of model capabilities and helps identify strengths and weaknesses based on community tests. We are releasing MERA to the research community to encourage innovation, establish standards, and promote reproducible evaluation practices.
@@ Acknowledgments
  • Contributors of RealCode, RealCodeJava, JavaTestGen: Pavel Zadorozhny, Adamenko Pavel, Rodion Levichev, Dmitrii Vorobev, Dmitrii Salikhov ,
@@
Appendix A Datasets examples
@@
ruHumanEval is the Russian counterpart of the HumanEval dataset Chen et al. (2021a), which assesses models’ abilities to generate solutions for straightforward programming problems in Python. The dataset contains the translated into Russian and manually verified tasks of the original dataset  [^15] https://huggingface.co/datasets/openai_humaneval, including the test cases, which were taken from Liu et al. (2023a) (10 test cases per task).
@@
Recent work has demonstrated that widely used surface-level metrics for code generation, such as BLEU Papineni et al. (2002b) and CodeBLEU Ren et al. (2020a), correlate poorly with functional correctness and real-world utility Chen et al. (2021b); Dong et al. (2025). These metrics tend to reward superficial similarity to reference implementations, often failing to capture whether the generated code satisfies the intended behavioral requirements Dong et al. (2025). Consequently, leading benchmarks such as HumanEval Chen et al. (2021b), MBPP Austin et al. (2021b), and AlphaCode (ClassEval) Li et al. (2022b) rely on execution-based evaluation, using pass@k (a task is considered solved if any of $k$ samples passes all tests) as the primary correctness metric. However, these benchmarks are typically limited to synthetic or isolated tasks, reducing their ability to reflect practical code completion and maintenance scenarios Zhuo et al. (2025).
@@
To overcome the “one-to-many” challenge in evaluation of the generation tasks, where multiple valid outputs exist for a single input, a novel evaluation methodology centered on an LLM-as-a-Judge was proposed. This framework leverages a large language model to assess the semantic equivalence between model-generated and human reference answers. Performance is reported using a pass@k metric, which credits a model if at least one of its $k$ generated candidates is deemed valid. The Qwen2.5-Coder-32B  [^16] https://huggingface.co/Qwen/Qwen2.5-Coder-32B model was used as the judge model, so its performance measurements on this benchmark may be biased. To verify the accuracy of the LLM-as-a-judge metric, comments were generated using GPT-4o for 45 examples from the dataset. These comments were annotated by human annotators for semantic equivalence with the original comments provided by real reviewers. Based on these annotated examples, the prompt for the Qwen2.5-Coder-32B judge model was refined until 100% agreement with human annotators was achieved on these 45 examples. Subsequently, the model’s agreement with annotators was further evaluated on an additional 100 comments generated in the same way using GPT-4o, achieving a Spearman correlation of 0.831 with human annotators and confirming its reliability for this task.
An example of RuCodeReviewer task:
@@
Next, methods and functions, along with their unit tests, were mapped using a method adopted from the paper Tufano et al. (2021). For Java, the mapping procedure was identical to the method by Tufano et al. (2021). For Python, Go, and C#, we develop similar mapping methods based on functions names and unique method invocation. For Javascript, the tests were mapped to functions by the last local method/function invocation in the test because test functions do not have identifiers when declared in it() and test().
@@
For the benchmark, the data was sampled for each language so as to cover different cases as evenly as possible in terms of the length of the focal function, the test function, and the entire context; and in terms of the type of both functions (function/method). The benchmark comprises a total of 2,500 samples, with 500 samples for each programming language. For evaluation, we use the CodeBLEU Ren et al. (2020b) metric to compare the original test from the repository and the generated test.
@@
CodeCorrectness is a novel benchmark designed to evaluate LLMs’ ability to identify whether a unit test is correct (compilable and executable) or incorrect (failing compilation or execution). Although advanced tools are being developed to assess generated code correctness Bui et al. (2025), dedicated benchmarks for evaluating LLMs’ capabilities in code assessment remain an underexplored research area.
@@
StRuCom Dziuba and Malykh (2025) addresses a critical challenge in AI-assisted code documentation by introducing the first large-scale dataset (153K examples) of structured Russian comments for Python, Java, JavaScript, C#, and Go. This resource enables robust evaluation of documentation generation models for Russian-speaking developers, filling the void of standardized benchmarks for non-English code explanations.
To construct StRuCom, the authors developed a hybrid sourcing methodology combining human-authored and AI-generated content. Human-written comments were extracted from over $150$K Russian GitHub repositories, identified through repository metadata analysis and license filtering. To ensure representation across all target languages, the corpus was supplemented with synthetic examples that were either created by Qwen2.5-Coder-32B Hui et al. (2024) or enhanced by Miqu-70B  [^17] https://huggingface.co/miqudev/miqu-1-70b. Every comment underwent automated validation against language-specific docstring conventions (GoogleDoc  [^18] https://google.github.io/styleguide/pyguide.html, JSDoc  [^19] https://jsdoc.app, etc.) using custom verification tools. The validation protocol enforced strict structural completeness, requiring all comments to comprehensively describe parameters, return values, exceptions, and types (if needed). For this benchmark, we curated a balanced subset comprising $7500$ comments for training ($1500$ per language) and $500$ for testing ($100$ per language).
For evaluation, the chrF metric Popović (2015) is employed due to its enhanced sensitivity to Russian morphological complexity. Unlike BLEU Papineni et al. (2002a) – which often fails to capture inflectional nuances – chrF operates through character n-gram weighting. This approach effectively detects subtle morphological shifts in case endings, verb conjugations, and derivational suffixes that are essential for assessing documentation fluency in Russian. The metric’s design aligns with the language’s rich inflectional system, providing a more reliable quality assessment for generated comments.
@@
The YABLoCo benchmark evaluates the proficiency of modern large language models (LLMs) in completing code functions based on textual descriptions from the user. While previous research has explored this capability for standalone functions Jiang et al. (2024), there has been limited investigation into functions intended for integration within larger code repositories. YABLoCo consists of 4 large repositories, with sizes ranging from 200,000 to 2,000,000 lines of code. The authors of the benchmark generated a call graph for each repository. For example, the LLVM repository contains approximately 175,000 functions and 571,000 inter-function calls. Next, the functions that do not correspond to the following criteria were filtered out from the benchmark: low test coverage, lack of developer comments, and excessively long function bodies. From the remaining candidates, the authors of YABLoCo manually selected 208 functions that featured clear docstrings aligned with their respective function bodies. These docstrings serve as the descriptive prompts for the LLMs being tested.
The main challenge for an LLM in generating a function for a specific code repository lies in the context size of the existing codebase. The new function must not only precisely follow the text description but also effectively utilize other functions and classes within the repository. This necessitates a careful selection of context that is both relevant for generation and not excessively lengthy. The authors of YABLoCo Valeev et al. (2025) discovered that an effective context for this purpose consists of the functions that are called by the target function. This context is referred to as the "oracle," as it is typically unavailable to the new function during generation, although it is known within the benchmark. Simply incorporating the oracle functions, without any modifications to the LLMs, resulted in an appreciable improvement on YABLoCo in the overall pass@10 metric, increasing it by 14 points (from 22.4 to 36.1).
@@
Appendix B Taxonomy details

### v3

@@ 4 Evaluation results and analysis
GPT-4.1 gpt4.1 - 0.381 0.382 0.144 0.418 0.416 0.443 0.450 0.344 0.555 0.096 0.201 0.297 0.660
Gemini 2.5 flash gemini2025 - 0.360 0.427 0.120 0.388 0.386 0.610 0.604 0.211 0.496 0.064 0.212 0.217 0.404
DeepSeek-Coder-V2-Inst deepseekcoderv2 236 0.350 0.360 0.149 0.364 0.386 0.433 0.392 0.269 0.494 0.060 0.188 0.200 0.714
GigaChat-2-Max mamedov-etal-2025-gigachat - 0.338 0.361 0.091 0.318 0.262 0.534 0.524 0.145 0.425 0.055 0.223 0.291 0.686
RuadaptQwen3-8B-Hybrid Tikhomirov_Chernyshov_2024 8 0.326 0.380 0.115 0.191 0.245 0.419 0.427 0.093 0.500 0.042 0.218 0.244 0.744
Qwen3-Coder-480B-A35B-Inst yang2025qwen3technicalreport 480 0.326 0.329 0.082 0.411 0.342 0.463 0.421 0.220 0.430 0.058 0.231 0.202 0.540
Kimi-K2-Inst-0905 kimiteam2025kimik2openagentic 1000 0.325 0.380 0.077 0.188 0.342 0.448 0.448 0.233 0.423 0.074 0.153 0.232 0.571
A-vibe HF Page 8 0.317 0.367 0.115 0.157 0.279 0.263 0.257 0.119 0.614 0.023 0.232 0.307 0.724
Qwen3-Coder-30B-A3B-Inst yang2025qwen3technicalreport 30 0.301 0.373 0.072 0.188 0.309 0.391 0.350 0.189 0.621 0.058 0.231 0.202 0.540
Qwen2.5-Coder-32B-Inst hui2024qwen25codertechnicalreport 32 0.300 0.306 0.111 0.323 0.386 0.311 0.289 0.220 0.466 0.065 0.206 0.213 0.519
GigaCode 1.4 Web page 32 0.293 0.166 0.135 0.322 0.352 0.357 0.305 0.313 0.027 0.064 0.230 0.276 0.676
Qwen2.5-72B-Inst hui2024qwen25codertechnicalreport 72 0.288 0.254 0.144 0.362 0.349 0.174 0.157 0.189 0.481 0.048 0.160 0.252 0.702
Seed-Coder-8B-Inst seed2025seedcoderletcodemodel 8 0.271 0.345 0.106 0.106 0.305 0.317 0.210 0.264 0.655 0.035 0.161 0.237 0.403
DeepSeek-V3.1 deepseekai2025deepseekv3technicalreport 685 0.248 0.259 0.106 0.084 0.275 0.345 0.282 0.189 0.230 0.065 0.159 0.293 0.372
GigaChat-2-Lite mamedov-etal-2025-gigachat - 0.229 0.269 0.048 0.099 0.148 0.309 0.251 0.035 0.316 0.025 0.107 0.144 0.712
Yi-Coder-9B-Chat ai2025yiopenfoundationmodels 9 0.206 0.181 0.135 0.067 0.255 0.350 0.173 0.229 0.145 0.016 0.168 0.192 0.364
Mixtral-8x22B-Inst jiang2024mixtralexperts 141 0.182 0.028 0.106 0.180 0.366 0.000 0.000 0.229 0.027 0.016 0.194 0.152 0.597
Kodify Nano Web page 1.5 0.181 0.226 0.082 0.041 0.104 0.191 0.173 0.053 0.235 0.003 0.112 0.092 0.475
CodeLlama70b-Inst roziere2023code 70 0.159 0.310 0.120 0.039 0.232 0.040 0.035 0.167 0.507 0.006 0.035 0.043 0.025
gpt-oss-120b openai2025gptoss120bgptoss20bmodel 120 0.155 0.106 0.096 0.350 0.309 0.004 0.012 0.273 0.208 0.000 0.107 0.055 0.093
DeepSeek-Coder-V2-Lite-Inst deepseekcoderv2 16 0.146 0.018 0.072 0.096 0.255 0.000 0.000 0.137 0.000 0.022 0.127 0.189 0.539
CodeLlama34b-Inst roziere2023code 34 0.141 0.097 0.120 0.019 0.195 0.000 0.000 0.088 0.144 0.023 0.176 0.120 0.472
CodeLlama7b-Inst roziere2023code 7 0.140 0.176 0.096 0.000 0.151 0.000 0.000 0.040 0.315 0.000 0.106 0.098 0.455
gpt-oss-20b openai2025gptoss120bgptoss20bmodel 20 0.095 0.019 0.038 0.305 0.285 0.000 0.000 0.013 0.044 0.001 0.148 0.062 0.115
@@
Table 3: MERA Code benchmark results with model specifications. The private tasks are CodeLinterEval, ruCodeEval and ruCodeReviewer. The board is captured on the 15.11.2025.
We evaluate both code-oriented and general-purpose models of various sizes. Table 3 shows the results of the models evaluation. Detailed results for public and private tasks are presented in the Appendix E. The snapshot of the leaderboard was taken on November 15, 2025.
Task complexity. The most difficult task is RuCodeReviewer which requires the largest number of skills according to our taxonomy. Moreover, the task presumes strong capabilities in Programming and Natural languages simultaneously, making it even harder. The easiest task is CodeCorrectness, probably due to its simple format (binary classification). In general, current benchmark results show modest model performance, with substantial headroom due to task difficulty.
Impact of natural languages. Since all prompts in MERA Code are formulated in Russian, a strong proficiency in the language is required for the models to get higher scores on some benchmark tasks. Hence, relatively small models, RuadaptQwen3-8B-Hybrid and A-vibe, targeted at the Russian language, appear at the top of the leaderboard. Particularly, A-vibe demonstrates superiority (0.307) on the StRuCom (generating commentaries in Russian) task.
Influence of the pretraining data. The Table 3 demonstrates that even smaller models may outperform larger ones from the same or “close” model families if they are trained specifically on the code tasks data. To be precise, Qwen2.5-Coder-32B-Instruct shows larger Total Score (0.3) compared to Qwen2.5-72B-Instruct (0.288). The same applies to the Qwen3-Coder-30B-A3B-Instruct result (0.301).
Model specialization. Our benchmark demonstrates that models lean toward certain specializations and programming languages. GPT-4.1 is efficient in Software Engineering, showing top performance on RealCode (0.418), RealCodeJava (0.416), JavaTestGen (0.344), and RuCodeReviewer (0.096) — the tasks that cover code completions, unit test generation, and code review. Gemini 2.5 flash achieves top scores on ruHumanEval (0.604) and ruCodeEval (0.610), which are designed for the evaluation of Python code generation abilities. DeepSeek-Coder-V2-Instruct is the best on the YABLoCo benchmark (0.149), outperforming other models in C++ and Long Context Comprehension. Seed-Coder-8b-Instruct is proficient in Python style guides (CodeLinterEval: 0.655). Such a distribution of skills among models confirms the complexity and diversity of our benchmark.
5 Conclusion.
@@
In this work, we present the MERA Code, a comprehensive evaluation tool designed for coding tasks and programming languages across eight different languages. MERA provides an open-source evaluation toolkit, a structured scoring system, and a public leaderboard, enabling standardized assessments of coding skills while addressing limitations in non-English contexts. This framework enhances our understanding of model capabilities and helps identify strengths and weaknesses based on community tests. We are releasing MERA to the research community to encourage innovation, establish standards, and promote reproducible evaluation practices.
@@ Acknowledgments
  • Contributors of RealCode, RealCodeJava, JavaTestGen: Pavel Zadorozhny, Pavel Adamenko, Rodion Levichev, Dmitrii Vorobev, Dmitrii Salikhov ,
@@
Appendix A Metrics description
Pass@k chen2021evaluating evaluates the functional correctness of the generated code.
Compile@k evaluates the correctness of the compilation of the generated code.
chrF popovic2015chrf is used in code-to-text tasks since its enhanced sensitivity to Russian morphological complexity.
BLEU papineni-etal-2002-bleu measures the n-grams level similarity of the prediction and gold answer.
CodeBLEU (CBLEU) ren2020codebleu combines regular BLEU with the measure of similarity of code syntax via abstract syntax trees (AST) and code semantics via data-flow.
Judge@k measures whether one of the top-k predictions matches with the gold answer via LLM-as-a-Judge zheng2023judging. The details about the validation of LLM-as-a-Judge could be found in Appendix C.
Appendix B Prompts language analysis
Model
YABLoCo
RealCode
RealCodeJava
JavaTestGen
CodeLinterEval
UnitTests
CodeCorrectness
pass@1 CBLEU EM
DeepSeek-Coder-V2-Instruct 0.077 0.054 0.003 -0.013 0.039 -0.028 0.001
Qwen3-Coder-30B-A3B-Instruct 0.0 -0.014 -0.094 -0.035 -0.006 0.007 -0.02
Qwen2.5-Coder-32B-Instruct 0.072 0.0 0.01 0.035 0.076 0.06 0.0
Qwen2.5-72B-Instruct 0.091 0.035 0.0 -0.035 -0.003 0.083 -0.009
GigaChat-2-Max 0.053 0.007 -0.07 -0.004 0.042 0.145 -0.002
GigaChat-2-Lite 0.01 0.04 0.013 -0.009 -0.023 0.02 0.002
A-vibe 0.058 -0.009 -0.04 -0.018 0.014 0.04 0.043
Kodify-Nano 0.058 -0.035 -0.044 -0.004 -0.035 -0.018 -0.011
Seed-Coder-8B-Instruct 0.067 -0.005 0.007 0.04 0.039 -0.06 0.014
Yi-Coder-9B-Chat 0.091 0.062 0.208 0.216 0.1 -0.14 0.032
Table 4: Detailed results on difference between scores with the Russian prompts and the English ones applied to the benchmark tasks. The difference is always “Russian — English”. Negative values mean that the score with the English prompts is larger than that with the Russian prompts.
We hypothesize that the language of the instructions does not affect the benchmark task scores. To test this, the original Russian prompts from MERA Code were translated into English using GPT-4.1, followed by LLM-as-a-Judge validation and manual corrections. We evaluated ten LLMs of varying scales, selected from a public leaderboard (see Table  3), encompassing both models specifically trained on Russian and general multilingual architectures. As we research the impact of the language of the task instructions, 4 out of 11 tasks (ruCodeEval, RuCodeReviewer, ruHumanEval, StRuCom) have been excluded: they contain natural language as part of either the input data or the golden answer, which does not fit into the hypotheses.
The Table 4 reports per-model and per-task score deltas between the Russian and English prompts (Russian – English). For most models and tasks, the differences remain close to zero (typically a few hundredths) and vary in sign, indicating no systematic advantage for either language. This presumably reflects the “English-default” nature of programming data: LLMs are widely trained on code paired with English documentation, comments, and instructions, so switching the instruction language often does not change the core code-centric signal and, therefore, has little effect on scores. The other explanation is that most tasks are dominated by the code and formal constraints, so models mainly focus on the programming content rather than the natural-language surface form. However, these two ideas are not contradictory.
Notable deviations are rare but informative. Yi-Coder-9B-Chat shows the largest gains from Russian on pass@1 tasks, especially RealCodeJava (+0.208) and JavaTestGen (+0.216), while UnitTests is one of the most mixed, with some models improving under Russian (e.g., GigaChat-2-Max +0.145) and others favoring English (e.g., Yi-Coder-9B-Chat -0.14). Overall, the benchmark appears robust to English-Russian prompt changes, with language effects emerging mainly for specific model-task combinations.
Appendix C Datasets examples
@@
ruHumanEval is the Russian counterpart of the HumanEval dataset chen2021evaluating, which assesses models’ abilities to generate solutions for straightforward programming problems in Python. The dataset contains the translated into Russian and manually verified tasks of the original dataset  [^11] https://huggingface.co/datasets/openai_humaneval, including the test cases, which were taken from liu2023your (10 test cases per task).
@@
Recent work has demonstrated that widely used surface-level metrics for code generation, such as BLEU 10.3115/1073083.1073135 and CodeBLEU ren2020codebleu, correlate poorly with functional correctness and real-world utility Chen2021EvaluatingLL; 10.1145/3695991. These metrics tend to reward superficial similarity to reference implementations, often failing to capture whether the generated code satisfies the intended behavioral requirements 10.1145/3695991. Consequently, leading benchmarks such as HumanEval Chen2021EvaluatingLL, MBPP Austin2021ProgramSW, and AlphaCode (ClassEval) Li2022CompetitionlevelCG rely on execution-based evaluation, using pass@k (a task is considered solved if any of $k$ samples passes all tests) as the primary correctness metric. However, these benchmarks are typically limited to synthetic or isolated tasks, reducing their ability to reflect practical code completion and maintenance scenarios zhuo2025bigcodebench.
@@
To overcome the “one-to-many” challenge in evaluation of the generation tasks, where multiple valid outputs exist for a single input, a novel evaluation methodology centered on an LLM-as-a-Judge was proposed. This framework leverages a large language model to assess the semantic equivalence between model-generated and human reference answers. Performance is reported using a pass@k metric, which credits a model if at least one of its $k$ generated candidates is deemed valid. The Qwen2.5-Coder-32B  [^12] https://huggingface.co/Qwen/Qwen2.5-Coder-32B model was used as the judge model, so its performance measurements on this benchmark may be biased.
LLM-Judge Validation To validate the automatic judge for the RuCodeReviewer task, we first generate candidate review comments with GPT-4o and manually annotate 300 generated–reference pairs. The goal of this step is to check whether the generated comment preserves the original semantic intent. Based on this annotated set, we refine the prompt for our LLM-based judge. After tuning, we test the chosen model Qwen2.5-Coder-32B on 100 Russian comment pairs. The model reaches a Spearman correlation of 0.831 with human ratings, which shows that it can reliably assess semantic equivalence.
For a broader check of consistency, we compare GPT-4o outputs using three different judges: Qwen2.5-Coder-32B, GPT-4o, and Claude-Sonnet-3.7. Their scores remain within each other’s standard deviation ranges, giving us confidence that Qwen2.5-Coder-32B is a solid choice. Since it is also faster and can be deployed locally, we adopt it as our default judge.
@@
RuCodeReview Judge Prompt
This prompt, translated from the original Russian formulation, is designed for an LLM-as-a-Judge:
You are a Judge.
You will be provided with a code change snippet and two code comments describing a problem. Your task is to determine whether both comments describe the same problem.
If the two comments describe the same problem and proposed solution, answer: correct. If the comments describe different problems and solutions, answer: wrong.
Write your answer, correct or wrong, without quotation marks or other extraneous characters.
Input data:
You’ll get the following data as input:
- Code block difference:
{diff_block}
- Comment 1: {comment1}
- Comment 2: {comment2}
Answer:
@@
Next, methods and functions, along with their unit tests, were mapped using a method adopted from the paper tufano2021unit. For Java, the mapping procedure was identical to the method by tufano2021unit. For Python, Go, and C#, we develop similar mapping methods based on functions names and unique method invocation. For Javascript, the tests were mapped to functions by the last local method/function invocation in the test because test functions do not have identifiers when declared in it() and test().
@@
For the benchmark, the data was sampled for each language so as to cover different cases as evenly as possible in terms of the length of the focal function, the test function, and the entire context; and in terms of the type of both functions (function/method). The benchmark comprises a total of 2,500 samples, with 500 samples for each programming language. For evaluation, we use the CodeBLEU ren2020codebleumethodautomaticevaluation metric to compare the original test from the repository and the generated test.
@@
CodeCorrectness is a novel benchmark designed to evaluate LLMs’ ability to identify whether a unit test is correct (compilable and executable) or incorrect (failing compilation or execution). Although advanced tools are being developed to assess generated code correctness bui2025correctness, dedicated benchmarks for evaluating LLMs’ capabilities in code assessment remain an underexplored research area.
@@
StRuCom dziuba2025strucom addresses a critical challenge in AI-assisted code documentation by introducing the first large-scale dataset (153K examples) of structured Russian comments for Python, Java, JavaScript, C#, and Go. This resource enables robust evaluation of documentation generation models for Russian-speaking developers, filling the void of standardized benchmarks for non-English code explanations.
To construct StRuCom, the authors developed a hybrid sourcing methodology combining human-authored and AI-generated content. Human-written comments were extracted from over 150150K Russian GitHub repositories, identified through repository metadata analysis and license filtering. To ensure representation across all target languages, the corpus was supplemented with synthetic examples that were either created by Qwen2.5-Coder-32B hui2024qwen2 or enhanced by Miqu-70B  [^13] https://huggingface.co/miqudev/miqu-1-70b. Every comment underwent automated validation against language-specific docstring conventions (GoogleDoc  [^14] https://google.github.io/styleguide/pyguide.html, JSDoc  [^15] https://jsdoc.app, etc.) using custom verification tools. The validation protocol enforced strict structural completeness, requiring all comments to comprehensively describe parameters, return values, exceptions, and types (if needed). For this benchmark, we curated a balanced subset comprising $7500$ comments for training ($1500$ per language) and $500$ for testing ($100$ per language).
For evaluation, the chrF metric popovic2015chrf is employed due to its enhanced sensitivity to Russian morphological complexity. Unlike BLEU papineni-etal-2002-bleu – which often fails to capture inflectional nuances – chrF operates through character n-gram weighting. This approach effectively detects subtle morphological shifts in case endings, verb conjugations, and derivational suffixes that are essential for assessing documentation fluency in Russian. The metric’s design aligns with the language’s rich inflectional system, providing a more reliable quality assessment for generated comments.
@@
The YABLoCo benchmark evaluates the proficiency of modern large language models (LLMs) in completing code functions based on textual descriptions from the user. While previous research has explored this capability for standalone functions jiang2024surveylargelanguagemodels, there has been limited investigation into functions intended for integration within larger code repositories. YABLoCo consists of 4 large repositories, with sizes ranging from 200,000 to 2,000,000 lines of code. The authors of the benchmark generated a call graph for each repository. For example, the LLVM repository contains approximately 175,000 functions and 571,000 inter-function calls. Next, the functions that do not correspond to the following criteria were filtered out from the benchmark: low test coverage, lack of developer comments, and excessively long function bodies. From the remaining candidates, the authors of YABLoCo manually selected 208 functions that featured clear docstrings aligned with their respective function bodies. These docstrings serve as the descriptive prompts for the LLMs being tested.
The main challenge for an LLM in generating a function for a specific code repository lies in the context size of the existing codebase. The new function must not only precisely follow the text description but also effectively utilize other functions and classes within the repository. This necessitates a careful selection of context that is both relevant for generation and not excessively lengthy. The authors of YABLoCo valeev2025yabloco discovered that an effective context for this purpose consists of the functions that are called by the target function. This context is referred to as the "oracle," as it is typically unavailable to the new function during generation, although it is known within the benchmark. Simply incorporating the oracle functions, without any modifications to the LLMs, resulted in an appreciable improvement on YABLoCo in the overall pass@10 metric, increasing it by 14 points (from 22.4 to 36.1).
@@
Appendix D Taxonomy details